OpenStudy (anonymous):

Anyone can help? Gloria has a lunch account in the school cafeteria. The balance, f(x), in her account can be represented by the function f(x) = −1.50x + 40, where x is the number of meals purchased. Suppose −1.50 changed to −2.75. What does this mean in the context of the problem? (2 points) Changing −1.50 to −2.75 means her initial balance has decreased. Changing −1.50 to −2.75 means her initial balance has increased. Changing −1.50 to −2.75 means the price per meal has decreased. Changing −1.50 to −2.75 means the price per meal has increased.

OpenStudy (nathaliama13):

B or D. :) I'd go for D because looks like she thought about buying more than one meal. :) so obviously the price has increased.

OpenStudy (anonymous):

Ty so much :)

OpenStudy (nathaliama13):

no problem!

OpenStudy (anonymous):

Can i confirm and see if this right the ones i did

OpenStudy (anonymous):

Ok So I got - Write the equation of a line with a slope of 4 and a y-intercept of −3. (2 points) 4x + y = −3 4x − y = −3 y = −3x + 4 y = 4x − 3 <My answer